International Workshop on Functional and (Constraint) Logic Programming (WFLP 2001)

Kiel, Germany, September 13-15, 2001

CALL FOR PAPERS

Program Committee Slim Abdennadher LMU München María Alpuente UP Valencia Sergio Antoy Portland State U Christoph Beierle Fern-U Hagen Alexander Bockmayr U Nancy Harold Boley DFKI Kaiserslautern François Bry LMU München Jürgen Dix U Koblenz Rachid Echahed IMAG Grenoble Uwe Egly TU Wien Moreno Falaschi U Udine Burkhard Freitag U Passau Thom Frühwirth LMU München Norbert E. Fuchs U Zürich Ulrich Geske GMD-FIRST Berlin Michael Hanus CAU Kiel (Chair) Herbert Kuchen U Münster Juan José Moreno Navarro UP Madrid Mario Rodríguez-Artalejo UC Madrid Dietmar Seipel U Würzburg

Local Organization Wolfgang Goerigk Michael Hanus Klaus Höppner Ulrike Pollakowski-Geuther Friedemann Simon Frank Steiner

Invited Speakers Peter Van Roy UC Louvain Thomas Arts Ericsson

Important Dates Submission: May 22, 2001 Notification: June 20, 2001 Final versions: July 20, 2001 Workshop: September 13 - 15, 2001 General: WFLP 2001 is the combination of two workshops of a successful series of annual workshops on declarative programming. The international workshops on functional and logic programming aimed at bringing together researchers interested in functional programming, logic programming, as well as their integration. Previous workshops have been held in Benicassim (Spain), Grenoble (France), Bad Honnef (Germany). The workshops on (constraint) logic programming are the annual meeting of the Society of Logic Programming (GLP e.V.) and bring together researchers interested in logic programming, constraint programming, and related areas like databases and artificial intelligence. Previous workshops have been held in Germany, Austria and Switzerland. In this year both workshops will be jointly organized in order to promote the cross-fertilizing exchange of ideas and experiences among researches and students from the different communities interested in the foundations, applications, and combinations of high-level, declarative programming languages and related areas. The technical program of the workshop will include invited talks, presentations of refereed papers and demo presentations.

Topics of Interest include (but are not limited to):

• Functional programming

• Logic programming

• Constraint programming

• Deductive databases

• Extensions of declarative languages, objects

• Multi-paradigm declarative programming

• Foundations, semantics, nonmonotonic reasoning, dynamics

• Parallelism, concurrency

• Program analysis, abstract interpretation

• Program transformation, partial evaluation, meta-programming • Specification, verification, declarative debugging

• Knowledge representation, machine learning

• Implementation of declarative languages

• Advanced programming environments and tools

• Applications

Submissions: Authors are invited to submit an extended abstract (no longer than 10 pages including figures and references) or a system description (no longer than 3 pages) in postscript format (11pt) before May 15, 2001. Submissions should include the title, authors’ names, affiliations, addresses, and e-mail. The details about the procedure to submit papers electronically are described on the workshop website. The primary focus is on new and original research results but submissions describing innovative products, prototypes under development or interesting experiments (e.g., benchmarks) are also encouraged. The proceedings with the full versions of all accepted contributions (no longer than 15 pages for regular contributions and 3 pages for system descriptions) will be published as a Technical Report of the University of Kiel. A special issue of ENTCS (Electronic Notes in Theoretical Computer Science) containing revised versions of high-quality contributions will be organized after the workshop. Best newcomer award: An award will be given to the best paper exclusively written by one or several young researchers who have not yet obtained their PhD degrees. Papers written in this category should be clearly marked “Student papers” in the submission.

             WFLP 2001 Web Page:  http://www.informatik.uni-kiel.de/~wflp2001/